OpenStudy (anonymous):

A triangular prism has a length of 8 inches. A larger similar triangular prism has a length of 12 inches. If 216 square inches of paper is used to wrap the smaller prism with minimum wastage, approximately how many square inches of paper will be required to wrap the larger prism? Answer 243 square inches 324 square inches 486 square inches 540 square inches

OpenStudy (dumbcow):

square the ratio of the sides 12/8 = 3/2 --> (3/2)^2 = 9/4 9/4 = x/216 solve for x

OpenStudy (anonymous):

would it be 324?

OpenStudy (dumbcow):

no, how do you get x by itself multiply both sides by 216 --> x = 216*(9/4)

OpenStudy (anonymous):

so 486 thank you